OpenStudy (anonymous):

|x - 3| + |x - 2| = 1

OpenStudy (anonymous):

x=2

OpenStudy (jiteshmeghwal9):

-(x-3)+(-1(x-2))=1 (x-3)+(x-2)=1 solve for x in both equations u will gt the possible values for x :)

OpenStudy (anonymous):

there are an infinite number of answers

OpenStudy (anonymous):

that is correct. you have to check the 3 regions, x<2, 2<x<3, and x>3. It turns out if 2<x<3, then it is a solution.

OpenStudy (anonymous):

x = 2 is valid, x = 2.1 is valid, x = 2.2 is valid, x = ... x = 3 is valid. any x between x = 2 and x = 3 works. question is how to arrive at this conclusion

OpenStudy (anonymous):

If x is greater than 2, then \[|x-2|=x-2\] if x is less than 3, then:\[|x-3|=-(x-3)=-x+3\]

OpenStudy (anonymous):

So for 2<x<3 we get:\[|x-2|+|x-3|=x-2-(x-3)=3-2=1\]

OpenStudy (jiteshmeghwal9):

well who thinks i'm incorrect ??

hartnn (hartnn):

not incorrect, but incomplete.

OpenStudy (anonymous):

there is a set of solutions. its the set [2,3]. any x such that:\[2\le x \le 3\]is a solution.

OpenStudy (anonymous):

knowing this is it possible to simplify \[\sqrt{x + 3 - 4\sqrt{x - 1}} + \sqrt{x + 8 - 6\sqrt{x - 1}} = 1\] using substitution ?

hartnn (hartnn):

if 'x' belong to Integer, only then there are 2 solutions, x=2,3

OpenStudy (jiteshmeghwal9):

This is the condition for absolute values :- When the value is unknown in the mod brackets we have to put the value in both positive & negative conditions. since |-x|=-(-x) & |x|=x

hartnn (hartnn):

\(x+3 -4 \sqrt{x-1} = (\sqrt{x-1})^2-2*2*\sqrt{x-1}+2^2\)

hartnn (hartnn):

=\([\sqrt{x-1}+2]^2\)

OpenStudy (anonymous):

How did you see that o.O

hartnn (hartnn):

because x+3 = x-1 +4 ...

OpenStudy (anonymous):

lol

hartnn (hartnn):

similarly tr for x+8-6 sqrt{x-1}

hartnn (hartnn):

hint : x+8 = x-1+9 :P

OpenStudy (anonymous):

very clever!

OpenStudy (anonymous):

[5, 10] ?

OpenStudy (anonymous):

Hmm... If all answers are between 2 and 3 are there still an infinite amount of answers or would the fact that everything outside of 2 and 3 mean it has to be finite?

hartnn (hartnn):

\(|\sqrt{x-1}-2|+|\sqrt{x-1}-3|=1 \\ y=\sqrt{x-1}\)

OpenStudy (anonymous):

if one circle is circumscribed about the other, does it contain more points because it has a greater circumference? (not really a question)

OpenStudy (anonymous):

thank you hartn, the solution was clear to me at your second post

hartnn (hartnn):

ok... circle circumscribed about other means ? circle inside a circle ?

OpenStudy (anonymous):

maybe the wrong terminology was used. consider one circle larger than the other but with the same center. does the outer circle contain more points than the inner because it has a greater circumference? the answer is no for a simple reason

hartnn (hartnn):

the outer circle will contain more points than the inner because it has a greater area.

OpenStudy (anonymous):

@binarymimic The question is are there infinite solutions to this problem? If so how is considered infinite but yet having the boundaries between 2 and 3...

hartnn (hartnn):

@patdistheman there are infinite numbers between 2 and 3.

OpenStudy (anonymous):

There is something intuitively that seems wrong about that. All the numbers fall between 2 and 3. We know there are numbers outside of the range then wouldn't by definition the answers not be infinite?

OpenStudy (anonymous):

Limitless or endless in space but also between 2 and 3 seem to be a contradiction.

OpenStudy (anonymous):

all of the numbers between 2 and 3 can be put in a 1 to 1 correspondence with the natural numbers, yes?

hartnn (hartnn):

infinite does not necessarily mean from -infinity to infinity , there are infinite numbers between 2 and 2.1 also.

OpenStudy (anonymous):

1 to 1 correspondence?

OpenStudy (anonymous):

for each number between 2 and 3 can you map it to a number in N ? where N = 1, 2, 3, ...

OpenStudy (anonymous):

Between 2 and 2.1 would have a limited supply no?

OpenStudy (anonymous):

perhaps not even a 1 to 1 correspondence with the natural numbers if we include irrationals, which are abundant.

OpenStudy (anonymous):

think of it this way. for every two positive numbers, their average is between them. and this number is also positive. you can do this forever if you like and continue to obtain numbers between 2.0 and 2.1

OpenStudy (anonymous):

It is clear that there appears to be an infinite supply of numbers between 2 and 3 that can be generated because additional numbers can always be added after but if that is infinite how would you describe the numbers outside of 2 and 3 while describing that case?

OpenStudy (anonymous):

i would say there are an infinite number of solutions, and an infinite number of non-solutions

OpenStudy (anonymous):

Putting a limit on infinite seems contradictory...

OpenStudy (anonymous):

its a paradox called continuity. it is kind of a problem but we tend to overlook it.

OpenStudy (anonymous):

Well I don't see the productive value at the moment. All I can say is it doesn't sit right.

OpenStudy (anonymous):

If numbers are used to describe something I imagine there is a limit at which the highest number of anything can be reached. Even all the atoms in the universe have a number...

OpenStudy (anonymous):

How would the difference between numbers that can relate to something and numbers that can not be defined...

OpenStudy (anonymous):

For example lets say there is 10^82 atom in the observable universe... How would you describe numbers beyond something that they can be applied to...

OpenStudy (anonymous):

well in mathematical space, how many points are there on the real line, for example? or on the plane, or in R3? etc. you reach a point at which you cannot define infinity by how many points it contains, per se, but by how the set containing it can be compared to other sets

OpenStudy (anonymous):

I suppose is essence the numbers have no point if they can not be associated to anything... I like the description... infinite solution and infinite non solution... I see no results on google... Was that just invented? lol

OpenStudy (anonymous):

you could argue that the region between 2 and 3 on the real line contains just as many points as R100 or R 10^83 or any finite value of n for Rn

OpenStudy (anonymous):

? i guess. i have no other way of describing it other than using interval notation.

OpenStudy (anonymous):

That doesn't sound correct either. I do not see why you would not be able to just keep adding zeros indefinitely. Yea I think we have something there. That may be a new form of description. How relevant that is to anything not sure.

OpenStudy (anonymous):

it is completely paradoxical to say that a part of a line contains just as many points as the entire line does, and also just as many points as a plane does, or all of 3D space, or all of any nth dimensional space.

OpenStudy (anonymous):

You heard it here folks... Openstudy.com... "Infinite Non Solution Numbers"

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I do not see how that is possible. I think the error is in thinking of mathematics as anything more than a tape measure.

OpenStudy (anonymous):

It would stand to reason that the distance between 1 to 2 is the same the distance between 2 to 3.

OpenStudy (anonymous):

You could have a tape measure with a length of 10^82 feet but what would you ever even do with it...